OpenStudy (anonymous):

WILL MEDAL OR FAN The area of a triangle is 323 square centimeters. The length of the base is 38 centimeters. What is the height of the triangle? @Directrix @blackops2luvr

OpenStudy (anonymous):

Hi @Agl202

OpenStudy (agl202):

Hey!

Nnesha (nnesha):

\[\huge\rm a = \frac{ 1 }{ 2 } b \times h\] replace variables by their values a = 323 b = 38 h =??

OpenStudy (anonymous):

@Agl202 do you know this one?

Nnesha (nnesha):

can you substitute a,b, values into the formula ??

OpenStudy (anonymous):

323 = 19 * h, now what?? @Nnesha

OpenStudy (anonymous):

are you gonna help me here or what?

Nnesha (nnesha):

19 is multiplying with h so do cancel out that form right side you need to do opposite of multiply which is divide so divide both side by 19

OpenStudy (anonymous):

oh, so h=17, you should of told me that in the first place.

OpenStudy (anonymous):

nevermind